OpenStudy (anonymous):

√(3+2√(2)) - √(3-2√(2)) I used a calculator to get the answer 2. How would I show that the calculated value is correct?

OpenStudy (amistre64):

by another calculator?

OpenStudy (amistre64):

you have a smaller number being subtracted by a bigger number, so that answer is gonna be +

OpenStudy (amistre64):

equate this this to some value "N" and solve it algebraiczlly

OpenStudy (amistre64):

typoed it :) you have a smaller number being subtracted FROM a bigger number, so that answer is gonna be +

OpenStudy (anonymous):

but how would I prove that the answer is 2, rather than any other positive integer?

OpenStudy (amistre64):

by solving for "N" of course. \(\sqrt{3+2\sqrt{2}}-\sqrt{3-2\sqrt{2}}=N\) what do you think we can do with this?

OpenStudy (amistre64):

or rather, how do we undo square roots?

OpenStudy (anonymous):

square both sides?

OpenStudy (amistre64):

id say thats a good start :) \(\left(\sqrt{3+2\sqrt{2}}-\sqrt{3-2\sqrt{2}}\right)^2=N^2\) now we can simplify this :)

OpenStudy (amistre64):

(a-b)^2 = (a^2 -2ab + b^2) if that helps

OpenStudy (amistre64):

\([3+2\sqrt{2}]-[2\sqrt{(3+2\sqrt{2})(3-2\sqrt{2})}]+[3-2\sqrt{2}]=n^2\)

OpenStudy (amistre64):

\([3+2\sqrt{2}]+[3-2\sqrt{2}]-[2\sqrt{(3+2\sqrt{2})(3-2\sqrt{2})}]=n^2\) \(6-2\sqrt{9-8}=n^2\) \(6-2\sqrt{1}=n^2\) \(6-2=n^2\) \(4=n^2=2\)

OpenStudy (amistre64):

well; N=2 it i can type :)

OpenStudy (anonymous):

Thank you so much :)